how do i solve:
6000x-500=y
- 9000x-950=y
and how do i find the x and y value

Answer by Cromlix(4381)

Rearrange:
6000x - y = 500
-9000x - y = 950
Subtract bottom from top
15000x = -450
Divide both sides by 15000
x = -450/15000
x = - 0.03
Substitute x = -0.03 into
6000x - y = 500
(6000*-0.03) - y = 500
-180 - y = 500
-y = 500 + 180
-y = 680
y = -680
x = -0.03 and y = -680
Hope this helps.
:-)